OpenStudy (anonymous):

multiply (x^5+y^4)(x^5-y^4)

OpenStudy (anonymous):

x^5*x^5-x^5*y^4+x^5-y^4-y^4*y^4 = x^5*x^5-y^4*y^4 = x^5+5-y^4+4 = x^10-y^8